Given the demand function for commodity x as qx=1500-0.25px2

solve the elasticity of demand for x if p= 20


1
Expert's answer
2021-05-17T10:37:54-0400

"q_x=1500-0.25(p_x)^2"

Slope of px

"\frac{dq_x}{d_x}= -0.5p_x=0.5(20)=-10"


"q_x=1500-0.25(p_x)^2"


"q_x=1500-0.25(20)^2"


"q_x=1500-0.25(400)"


"q_x=1500-100"


"q_x=1500-100"


"Price\space elasticity\space of \space demand=Slope \space of\space P \times \frac{p}{q}"


"E=-10\times \frac{20}{1400}"


"E=-0.14"